OpenStudy (anonymous):

What is the simplified different quotient of f(x)=x^2+4x-3?

OpenStudy (anonymous):

2x+4

OpenStudy (anonymous):

My book has the answer as 2x+h+4, was wondering if you could please go step by step with me you obviously know the correct steps just left out something, thanks

OpenStudy (anonymous):

simplified different quotient of f(x) is, =[f(x+h)-f(x)]/h here f(x)=x^2+4x-3 so, s.d.q=[{(x+h)^2+4(x+h)-3}-{x^2+4x-3}]/h =[x^2+2xh+h^2+4x+4h-3-x^2-4x+3]/h =[2xh+h^2+4h]/h =2x+h+4